OpenStudy (skittles_for_life6422):

WILL GIVE MEDAL AND FAN!!! What is the median of the data set below? round to the nearest whole number.

OpenStudy (anonymous):

What's your question?

OpenStudy (skittles_for_life6422):

\[SIMPLIFY. \left[\begin{matrix}4 & 5 \\ 0 & -1\end{matrix}\right] +2\left[\begin{matrix}1 & 8 \\ -4 & -6\end{matrix}\right]\]

OpenStudy (skittles_for_life6422):

How do you solve it

OpenStudy (anonymous):

I have no idea....

OpenStudy (anonymous):

first multiply the 2 into the second matrix

OpenStudy (skittles_for_life6422):

ok

OpenStudy (anonymous):